1- Suppose a firm earns Net Income of $1,200,000. The company pays an Ordinary Dividend of $400,000 and a Preference Dividend of $200,000. Throughout the financial year, the firm has 100,000 Ordinary Shares and 200,000 Preference Shares. The firm’s Earnings Per Share (EPS) is:

2- Suppose a firm earns Net Income of $1,000,000. The company does not pay dividends.

At the start of the financial year the firm had 980,000 Ordinary Shares. On 31 March, the firm issued 20,000 Ordinary Shares for Cash. On 30 September, the firm issued a further 20,000 Ordinary Shares for Cash.  The firm’s Earnings Per Share (EPS) is:

3-

Suppose a firm issues 2,000 Convertible Preference shares with a par value of $2 per share. The shares were issued at a price of $50 per share. The journal entry to record the transaction at the date of issue is:

a. Debit: Cash $100,000. Credit: Share Capital - Preference, $96,000. Share Premium – Conversion Equity: $4,000.

b. Debit: Cash $100,000. Credit: Share Capital - Preference, $4,000. Share Premium – Conversion Equity: $96,000.

c. Debit: Cash $100,000. Credit: Share Capital - Preference, $4,000. Share Premium – Preference: $96,000.

d. Debit: Cash $200,000. Credit: Share Capital - Preference, $199,000. Share Premium – Preference: $1,000.

e. None of these answers
Answer #1

1.
Net Income attributable to common stockholders = Net Income - Preference Dividend
= $1200000 -200000 = $1,000,000

EPS = Net Income attributable to common stockholders / Ordinary Shares
= $1,000,000 / 100000 = $10 per share

2.
Weighted Average Stock outstanding = 980000 + 20000 x 9/12 + 20000 x 3/12 = 1000000
EPS = $1000000/1000000 = $1 per share

As per HOMEWORKLIB RULES, we are suppose to answer 1 question, i have answered 2, so kindly post others separately